40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 2 07:54:55 2010, steve, Update, MOPA, laser headtemp is up htemp.jpghtemp20d.jpg
    Reply  Wed Jun 2 11:28:31 2010, Koji, Update, MOPA, laser headtemp is up 
Message ID: 3033     Entry time: Wed Jun 2 07:54:55 2010     Reply to this: 3035
Author: steve 
Type: Update 
Category: MOPA 
Subject: laser headtemp is up 

Is the cooling line clogged? The chiller temp is 21C See 1 and 20 days plots

Attachment 1: htemp.jpg  68 kB  | Show | Show all
Attachment 2: htemp20d.jpg  75 kB  Uploaded Wed Jun 2 08:59:20 2010  | Show | Show all
ELOG V3.1.3-